RAID Project Updates – October/November

The RAID Scientific Drilling Project Team continues to send updates as the project progresses.  Here are the updates for late October and November 2016.
See December updates here.

Check back here for regular updates as they are added. See the original project description here. 

OCTOBER 2016

10-21-16

Chris Delahunty, Rich Szentmiklosi, Mike Vinson, and AJ Vecchiarelli all depart for Christchurch, New Zealand.

10-24-16

raid-scientific-drilling-img_20161024_120616

10-25-16

Raid team was delayed yesterday from departing CHCH to MCM by airplane and weather issues. We received a call this morning about persistent weather issues and have been delayed another 24 hrs. We are currently in standby.

raid-scientific-drilling-img_20161025_094216

10-26-16

We are loaded onto a C-17 and on our way to MCM.  Two days late but moving now. Mike Vinson is excited.

10-27-16

We are successfully into McM today.  We had a very nice ride on the C-17 (much better than the C-130s).  We are headed to training today.  I will get the drill logs out of the container today and start the drill logs.

raid-scientific-drilling-img_20161027_143028

raid-scientific-drilling-img_20161027_085321

10-28-16

We have successfully completed the science brief, vehicle training, fire, waste, medical and Pisten Bulley training.  Still another training day today.

10-29-16

Finished with almost all training.  We have vehicle walk around Monday and small snow machines Tuesday.  We have been into the containers today and are working a game plane on lifting the containers to the skis.  We will be working with the USAP crew on Saturday and Monday to get the skis assembled and then we will use the big forklift to set the FRS, POWER, and SHOP containers on skis as Tuesday/Wednesday, then move onto the drill and rod rack Thursday/ Friday.raid-scientific-drilling-img_20161029_142211

10-31-16

Today we finally got some traction here.  With the help of the heavy machine ops group we were able to move all the DES cargo to the workpad near the dock and all of the skis and miscellaneous parts for the skis.  While disassembling the ski packs and flipping them over, we had a problem where one of the plastic sheets on the bottom of the ski was fork lift torn and we needed to replace the sheet.  That was an unexpected issue.  The ski units were not previously assembled, so we started that this afternoon.  We have one full set of skis completed with another set left to do tomorrow morning.  The crane that we need to lift the containers is being repaired and so we will use a forklift to do the picks of the FRS, POWER, and SHOP onto their respective skis tomorrow afternoon or Wednesday morning.   We will be stripping off the winter covers from the drill and rod rack tomorrow as well. I forgot to take a picture today so I will get one for tomorrow.  Looking at our schedule we are behind in that we were delayed 2 days in CHCH, we did our training on Friday and Saturday, then we were ready to work yesterday and McM was pretty much on Sunday shutdown.  With us doing the ski assemblies we will need Tuesday morning to finish the Shop and Power skis and then Wednesday all day to do the Drill and Rod skis.

NOVEMBER 2016

11-1-16

Today we were able to set the FRS on skis and move the shop and power containers to the staging area.  We set up 1 ski for the rod rack.  And we are waiting for a crane to reclaim 2 skis from the wissard equipment. Tomorrow we will break into the Rod and drill racks to remove the on board storage.  I have attached some pics from the days’ ops.

raid-scientific-drilling-img_20161101_162300

raid-scientific-drilling-img_20161101_165546

11-2-16

We were able to get the last 2 needed skis from the wissard equipment today.  The crane lifted off a container on them.  This is great news that we now have a working crane.  We lifted the ROD skid and took off its winter cover.  Very minimal snow drifting past the cover and none on the rod skid deck.  We off loaded all packed items on the rod skid.   We can now finish the outfit of the POWER and FRS containers.  We also completed all assembly of the drill and rod skis today.  Looking forward to all containers on skis by Friday.

raid-scientific-drilling-img_20161102_190446

raid-scientific-drilling-img_20161102_190416

raid-scientific-drilling-img_20161102_190453

11-4-16

So for today and yesterday, we have the Power container up on skis, the FRS has been taken to Willey field. We finished the drill and rod ski sets. The drill is un-tarped. The drill and rod are ready to lift and put on their skis.  The shop is ready to put on its skis.  So tomorrow will be a big day to get all equipment on skis.

We have had to do quite a bit of fabrication over the last 2 days to get the old ISO 2 ski sets retrofitted to make up to the new ISO 3 walking beam skis.

raid-scientific-drilling-img_20161104_140652

11-5-16

We have made much progress over the last 2 days.  All containers are on their respective skis. The jack legs are all assembled on the drill and rod skids.  The Power container is externally completed with the damper housing, exhaust lines, cable bridge, safety fall restraint, and ladders.  It is looking very sharp.  The Shop container is complete with cable bridge.  The drill rod was loaded down to sausage point. And we installed three of the J brackets to the skis and drill today. We will probably move the power and shop skid to Willey field tomorrow and possibly the rod skid. When the components are to Willey field we will then start the interior assembly getting all the DNF items placed.

raid-scientific-drilling-img_20161105_074152

raid-scientific-drilling-img_20161105_130437

raid-scientific-drilling-img_20161105_130443

raid-scientific-drilling-img_20161105_130537

raid-scientific-drilling-img_20161105_130601

raid-scientific-drilling-img_20161105_165658

11-7-16

Another good day.  We were able to drill and place all the J clips into the ISO 3 Ski benches for the drill and rod skids.  We assembled the catwalks to about 50%.  We were scrambling to get all the loose items loaded into the rod skid because we thought that we would be moving the drill and rod skids to Willey field tomorrow.  Because of lack of snow the fleet ops group has been stockpiling snow along the road from sausage point to the town side transition to the ice.  They aren’t quite as far as they would like to be so it looks as though the move day for the drill and rod will be Wednesday. We did move the Power and shop out today from the wharf.  I have put video and pics of today’s endeavors on the google drive. Tomorrow and Wednesday we will be working on the interior of the skids to start placing all the DNF items.

raid-scientific-drilling-img_20161106_152740

raid-scientific-drilling-img_20161107_111832

DCIM105GOPRO

11-8-16

We completed as much of the Drill and Rod racks as we could before towing out to Willey field.  In order to get to the sea ice, the fleet ops needed to get about a ¼ of a mile worth of road paved with snow.  This area is normally always snow packed, but as with everything RAID it seems we hit a low snow year with higher than average temperatures.  This has been nice for the assembly of the units, but now we see a minor downside.  The snow was trucked in from point McM where some drifts have occurred, and is ready for the bulldozers to flatten and pack it.  We have loaded all RAID DNF boxes and equipment to be hauled out Willey field and as soon as the drill and rod racks can be towed out we will resume set up at that point.  We are currently waiting on delivery and containers to resume.

raid-scientific-drilling-pano_20161109_160544

11-9-16

We have successfully moved all modules out of McM to the SpoTsa site.  It was a blustery cold day here and it took a lot of help from the fleet ops group to get the snow road together as well as the tremendous help with forklifts, cranes and man labor.  If you have time check out the videos especially the one labeled GOPR Tractor pull.raid-scientific-drilling-img_20161109_074122

11-11-16

Today we were able to get the Power Module completely finished with respects to rigging up out of long term storage.  We are making good progress every day.  Our goal is to have the FRS done tomorrow.  The USAP Electricians we onsite today running electrical in the power module for the galley and tractor plugin stations. Things are definitely moving ahead.

11-12-16

Today we started out with the hopes of refitting the FRS module with all of its DNF items, but we got word from the traverse group that they wanted to move the containers so that they were in operational position.  We changed gears and started laying out the rig mats. Then, after lunch, the tractors started pulling containers into location.  By 16:30 we had the modules in their relative locations. Then, with the time remaining, we worked on the Drill module DNF items and are about 50% complete.

11-14-16

We are 50% complete with DNF in the FRS and the cable bridges are in place.  We need to move the drill closer to the Power module in the future in order to get that cable bridge to work better.  For now the lines are on the ground.

11-15-16

We have finished all the work we can do prior to having our generators started and functioning.  We are about 92% overall complete, with the remaining 8% in FRS Glycol load into the boiler, generator startup, Power check at all levels of each container.  And final leveling, joining and raising of the drill mast and tent. Rich and Chris will be in an all-day crevasse training for those going on the traverse.  We hope the technician will be available Thursday day to finalize the generators.  We are set to traverse next Wednesday.

11-16-16

We have one generator running as of 1700 hrs today! We did a power check on the power module and the shop module.  They are 100% checked off.  We are working to tweak the damper controls, and were able to get them to work pretty well today holding the interior temperature around 60 – 70 F.  We finished the junk baskets on the Shop Container.  Rich and Chris were in crevasse training from 0800 to 1700 today.  Climbing a rope and descending the rope is easily accomplished by Rich, he has some funny video of Chris taking a half hour to do what he can do in 10 minutes.  The outside of the modules were wiped down to get rid of the McMurdo crud caked on them.  The drill doors/rod doors latches are complete. The drill shipping floor removed. Tomorrow we are going to try to marry the rod and drill modules and maybe have the tent up.

11-17-16

We have generator B running today and complete control over the Power Module.  Generator A ran all day and kept good pace with us.  We had all of the modules powered up and running lighting plugs and heaters today and with the drill running while we were marrying the drill and rod modules, there was only a 20% load on the system.  We are ready to step the mast and auger a hole.

11-18-16

The tent is up! We have our first picture with the entire operation assembled. We had distinguished visitors out today in Scott Borg, Thomas Wilch, and Nature McGinn.  The day started of snowy with light winds.  We were debating on whether to roll back the roofs and set up the tent as we had gotten word that Jessie Crane had left the continent, when Matt Kippenhan said he had a bunch of people that wanted to see it.  So we put up the tent in 2 hours and the wind was blowing, and snow everywhere.   But we did it. We are in the 99th percentile complete.  Just a few more things to check off.  We will drill an auger hole tomorrow at 1300.

11-19-16

We were able to auger a hole today to the depth of 30′ below the surface (fbs).  After the augers were pulled we measures 22′ of open hole.  That is 78% cleared.  Better than the 50% experienced in January, but we still need to refine the removal process.  We continued today to do odd jobs, making the layout better, checking things off of our lists.  We will be traversing on Wednesday the 23rd, and in order to do that we will need to be completely stowed by the 21 EOD as the traverse crew wants to be off on Tuesday to get personal things done. So we will be stowing things tomorrow in prep for traverse.

11-20/21-16

We successfully broke down the entire drill platform system in about 11 hours to the point we were ready to traverse. This occurred over Sunday and Monday 20-21.  We can expect some efficiency to come from doing this a few more times, but we were able to complete the task in fairly inclement weather over the last 2 days.  We are currently staged at SPoTSa ready for a demob of Wednesday the 23rd.  Standby, WOT: Waiting on Traverse.

11-22-16

We are personally prepping for the mobilization tomorrow.  Currently on Standby: Waiting on Traverse.

11-23-16

We are leaving today around 1100 hrs.  The traverse train will be 5 tractors.  I will take pictures as I can.  I will send daily updates as I can, but you might not hear from us for several days.

11-24-16

Traverse

11-25-16

We have arrived!  We made it to the MB-A1 site @ 1100 hrs today and started Track Packing the pad.  I have attached the logs from the last 3 days and a photo.

11-26-16

We were able to track pack the site yesterday and let it sinter overnight.   We are ready to start moving equipment onto the pad now.

11-27-16

We have successfully moved the Drill, Rod, Power, FRS and Shop Modules onto the drill pad at Hole A1.1.  The Traverse team moved their modules out and set their camp up. The Camp Manager and her team will fly in tomorrow and the remainder of the camp will be prepped.  We anticipate John, Jeff, Blaise and the other drill team will be on site around Dec. 1.

11-28-16

I had my hopes set on having the tent up today but we had to do some repairs on a few seams that were unraveling. Those are accomplished now and tomorrow looks good.

11-29-2016

The tent is up and we have checked the FRS boiler.  We are ready to do some augering today!  In attached picture, note the white speck above the drill.  That is a LDB balloon.

11-30-16

We were able to auger a hole to 154.75 fbs yesterday.   We ran the video bore scope and found fill at the 124 fbs mark that is roughly 20% still left in the hole, that is too much.   So we are working a plan to evacuate the remaining chips out.

See December 2016 updates here.